OpenStudy (anonymous):

Statistics.. Grace loves to eat salad! How many salads can she put together if she can pick out one type of lettuce from 2 choices, one vegetable from 4 choices and one dressing from 7 choices?

OpenStudy (shane_b):

2*4*7=?

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

It's that simple?

OpenStudy (shane_b):

Yes...since she can only pick one of each and the order of her choices doesn't matter.

OpenStudy (anonymous):

oh ok.. Do you have time for one more?

OpenStudy (shane_b):

I do...but prob & stats isn't my thing. This was just an easy one :)

OpenStudy (anonymous):

its kinda the same type of question. A restaurant offers four sizes of pizza, two types of crust, and eight toppings. How many possible combinations of pizza with one topping are there?

OpenStudy (shane_b):

Same thing...since she's only picking one of each item and again, the order doesn't matter. You may also see it calculated as a product of ratios:\[\frac{1}{4}*\frac{1}{2}*\frac{1}{8}=\frac{1}{64}\]So there are 64 difference combinations available.

OpenStudy (shane_b):

err *different

OpenStudy (anonymous):

Okay Great. So there was another problem that asked PA license plates have 3 letter followed by 4 number. If the same letter cannot be repeated, how many can be made?